# Align-to-Distill: Trainable Attention Alignment for Knowledge Distillation in Neural Machine Translation

### 03 Mar 2024 v1

@@ Abstract
The advent of scalable deep models and large datasets has improved the performance of Neural Machine Translation (NMT). Knowledge Distillation (KD) enhances efficiency by transferring knowledge from a teacher model to a more compact student model. However, KD approaches to Transformer architecture often rely on heuristics, particularly when deciding which teacher layers to distill from. In this paper, we introduce the “Align-to-Distill” (A2D) strategy, designed to address the feature mapping problem by adaptively aligning student attention heads with their teacher counterparts during training. The Attention Alignment Module (AAM) in A2D performs a dense head-by-head comparison between student and teacher attention heads across layers, turning the combinatorial mapping heuristics into a learning problem. Our experiments show the efficacy of A2D, demonstrating gains of up to +3.61 and +0.63 BLEU points for WMT-2022 De→Dsb and WMT-2014 En→De, respectively, compared to Transformer baselines.
@@
Heegon Jin, Seonil Son, Jemin Park, Youngseok Kim, Hyungjong Noh, Yeonsoo Lee NC Research, AI Tech Center {heegon, deftson, hbk5844, kys159, nohhj0209, yeonsoo}@ncsoft.com
@@ 1 Introduction
In this paper, we introduce a novel KD strategy, Align-to-Distill (A2D), that addresses the feature mapping problem using a trainable Attention Alignment Module (AAM). Unlike earlier KD methods that relied on combinatorial feature mapping heuristics, A2D provides an end-to-end trainable solution. This advancement is facilitated by the AAM, which aligns student attention maps with those of the teacher, resulting in more effective distillation. The adaptive alignment of features also removes the necessity for a data-dependent mapping strategy. AAM enables each attention head in the student model to be compared with every head in the teacher model across different layers, by employing pointwise convolution with only a few additional parameters. As a result, there is no longer a need for head or layer parity between the student and teacher models.
@@ 3 Related Works
Figure 1: Attention Transfer with A2D. The Attention Alignment Module (AAM), implemented as a pointwise convolution layer, produces intermediate attention maps from a collection of student attention maps. The number of intermediate maps matches the total attention maps of the teacher model, encompassing all layers and heads. These intermediate maps are then directly compared to the teacher’s attention maps using KL-Divergence, without any form of reduction. The same will occur for every layer of the student model, facilitating dense knowledge distillation.
@@ 5.4 Results on Low-resource datasets
Low-resource datasets present a unique challenge for NMT models, emphasizing the importance of effective knowledge distillation. Table 1 summarizes all the results for low-resource NMT. Models are compared with BLEU scores computed using sacreBLEU Post (2018), with a confidence interval of 95%, and the number of bootstrap resamples is 1000. Previous KD methods, such as TinyBERT and MiniLM, impose a constraint that the number of heads in the student must match that of the teacher. In contrast, A2D does not have this limitation regarding the number of attention heads, allowing us to train students with 8 heads using teachers with only 4 heads. We also present the results of students with 8 heads in section 6.

### 25 Mar 2024 v3

@@ Abstract
The advent of scalable deep models and large datasets has improved the performance of Neural Machine Translation (NMT). Knowledge Distillation (KD) enhances efficiency by transferring knowledge from a teacher model to a more compact student model. However, KD approaches to Transformer architecture often rely on heuristics, particularly when deciding which teacher layers to distill from. In this paper, we introduce the “Align-to-Distill” (A2D) strategy, designed to address the feature mapping problem by adaptively aligning student attention heads with their teacher counterparts during training. The Attention Alignment Module (AAM) in A2D performs a dense head-by-head comparison between student and teacher attention heads across layers, turning the combinatorial mapping heuristics into a learning problem. Our experiments show the efficacy of A2D, demonstrating gains of up to +3.61 and +0.63 BLEU points for WMT-2022 De→Dsb and WMT-2014 En→De, respectively, compared to Transformer baselines.  [^1] The code and data are available at https://github.com/ncsoft/Align-to-Distill.
@@
Heegon Jin, Seonil Son, Jemin Park, Youngseok Kim, Hyungjong Noh, Yeonsoo Lee NC Research, AI Tech Center Seongnam, South Korea {heegon, deftson, hbk5844, kys159, nohhj0209, yeonsoo}@ncsoft.com
@@ 1 Introduction
In this paper, we introduce a novel KD strategy, Align-to-Distill (A2D), that addresses the feature mapping problem using a trainable Attention Alignment Module (AAM). Unlike earlier KD methods that relied on combinatorial feature mapping heuristics, A2D provides an end-to-end trainable solution. The adaptive alignment of features removes the necessity for a data-dependent mapping strategy. Furthermore, AAM aligns the student attention map in each head with those of the teacher, resulting in more effective distillation compared to layer-wise feature mapping. AAM enables each attention head in the student model to be compared with every head in the teacher model across different layers, by employing pointwise convolution with only a few additional parameters. As a result, there is no longer a need for head or layer parity between the student and teacher models.
@@ 3 Related Works
Figure 1: Attention Transfer with A2D. The Attention Alignment Module (AAM), implemented as a pointwise convolution layer, produces intermediate attention maps from a collection of student attention maps. The number of intermediate maps matches the total attention maps of the teacher model, encompassing all layers and heads. These intermediate maps are then directly compared to the teacher’s attention maps using KL-Divergence, without any form of reduction.
@@ 4.1 Attention Map as a Knowledge
The attention mechanism is a vital component in Transformer models, capturing the context and relationships between words in a sentence. Attention alignment ensures that the student model, despite potential differences in network structure compared to the teacher model, focuses on the same word relationships and contextual nuances as the teacher. This process allows the student model to gain insights beyond mere output mimicry; it learns the underlying relationships that the teacher model uses for prediction, effectively transferring knowledge from the teacher model to the student.
@@ 5.4 Results on Low-resource datasets
Low-resource datasets present a unique challenge for NMT models, emphasizing the importance of effective knowledge distillation. Table 1 summarizes all the results for low-resource NMT. Models are compared with BLEU scores computed using sacreBLEU Post (2018), with a confidence interval of 95%, and the number of bootstrap resamples is 1000. Previous KD methods, such as TinyBERT and MiniLM, impose a constraint that the number of heads in the student must match that of the teacher. In contrast, A2D does not have this limitation regarding the number of attention heads, allowing us to train students with 8 heads using teachers with only 4 heads. We also present the results of students with 8 heads in section 6. Doubling the number of attention heads from 4 to 8 does not alter the overall parameter count in the ’A2D (8heads)’ model. This is achieved by proportionally reducing the dimension per head ($d_{head}=d_{model}/n_{head}$) while keeping the feature dimension ($d_{model}$) constant, thus ensuring a fair comparison.
@@
In low-resource settings, a student model could outperform the teacher under the guidance of the teacher model. This is due to the enhanced generalization provided by the regularization effect of distillation, as discussed in Mobahi et al. (2020); Yuan et al. (2020). The teacher model, often being larger and more complex, captures rich feature representations (such as attention maps) of the data. Through distillation, the student model learns these representations, which might be more generalizable than those learned from the relatively small amount of raw data alone.
@@ 7 Conclusion
8 Limitations and Future work
Although A2D is architecturally flexible without constraints on hidden size or number of attention heads, the teacher and student models must share the same vocabulary. This requirement potentially restricts its broader applicability. Also, while our work demonstrates A2D’s effectiveness on the decoder module in Table 5, we have not yet tested A2D on decoder-only models. The scope of this paper primarily focuses on encoder-decoder-based translation models. In future work, we plan to extend A2D’s application to decoder-only models. Moreover, the concept of A2D may encompass a broader range of architectures in future work, as the idea of adaptively aligning features is not restricted to using attention as a feature.